A dielectric of dielectric constant 2.5 is filled in the gap between the plates of a
capacitor. Calculate the factor by which the capacitance is increased, if the
dielectric is only sufficient to fill up one-fourth of the gap
1
Expert's answer
2020-03-04T10:01:15-0500
"C=\frac{\epsilon_0 A}{d}"

We can consider the capacitor to be series connection of two capacitors. 


"\frac{1}{C'}=\frac{1}{C_1}+\frac{1}{C_2}"

"C_1=\frac{\epsilon_0 A}{0.75d}=\frac{4}{3}C"

"C_2=\frac{\epsilon\epsilon_0 A}{0.25d}=\frac{2.5\epsilon_0 A}{0.25d}=10C"

"\frac{1}{C'}=\frac{3}{4C}+\frac{1}{10C}"

"C'=\frac{20}{17}C"

The capacitance is increased by 


"\frac{20}{17}\approx 1.18"